# Not Every Gift Comes in Gold Paper or with a Red Ribbon: Exploring Color Perception in Text-to-Image Models

### v1

@@ Abstract
{strip}[Uncaptioned image]
Figure 1: Do text-to-image models such as FLUX [11] (top row) and Stable-Diffusion-2.1 [32] (bottom row) accurately perceive colors? As shown above, modern models can faithfully render even uncommon colors—e.g., cornflower-blue—in simple, single-object prompts. However, when faced with multi-color, multi-object prompts, their performance degrades significantly. To evaluate this gap, we contribute a dedicated benchmark, along with an editing method that corrects these semantic errors. Our method consistently outperforms existing editing approaches, such as AnySD [36], FPE [18], and MasaCtrl [6], which struggle to resolve such misalignments.
@@ 1 Introduction
Unfortunately, the answer might be no. Evidence suggests that these models struggle in faithfully generating the semantics conveyed by complex multi-subject prompts [8, 29, 9]. Consider the results depicted in Figure 1. While text-to-image models mostly succeed in generating a “cornflower-blue gift box”, they are challenged by the more complicated prompt that also contains additional colored entities. We observe that a case study on colors offers a unique opportunity for performing a rigorous evaluation over text-to-image generative models. By contrast to coarse CLIP-based metrics, entangled VQA-based scores [17], or small-scale human evaluations conducted by prior work, performance can be measured explicitly using perceptual color differences.
Our contribution is twofold. First, we conduct a set of controlled experiments, comparing results generated using a single object (and its associated color) and results generated using multiple objects (and multiple colors). We also separately consider perceptually-close colors and distant colors, and contribute a benchmark for color understanding in image generation. Second, based on insights from this analysis, we develop a flexible and model-agnostic approach for editing images to precisely match multiple color specifications while preserving all other attributes (1).
Our analysis reveals that pretrained text-to-image models are significantly more successful in representing color attributes in the single subject setting. Inference-time methods that explicitly target the problem of multi-attribute leakage cannot robustly bridge this gap, and also often compromise performance over single subject prompts. We also observe a clear performance gap between close and distant color pairs, demonstrating that in most cases distant color pairs yield worse performance, due to possible leakage which more severely affects performance.
@@ 2 Related Work
Prior work [8, 29] has analyzed and identified typical misalignments, such as catastrophic neglect (i.e., failure to generate one or more mentioned subjects) and incorrect attribute binding (i.e., failure to match between attributes and subjects). Various mitigation schemes have been proposed. These increase semantic alignment by modifying the text embeddings [12, 34] or by better aligning the cross-attention maps [8, 29]. Several techniques mitigating semantic misalignments also utilize additional inputs, such as fine-grained segmentation maps [16] or coarse spatial layouts [28, 9]. In this work, we focus on the problem of incorrect attribute binding, focusing specifically on colors. We introduce a new benchmark for studying such semantic misalignments, analyzing the performance of existing text-to-image models on our proposed benchmark.
@@
Another line of work has explored color understanding for comparing multimodally-trained models with text-only models. Alper et al. [3] demonstrated that vision-and-language models can better associate objects with colors, in comparison to text only models, like BERT [10]. Several work [24, 19] have studied the effect of reporting bias on language models, demonstrating that they are able to connect object and colors despite the reporting bias in text [22]. Monroe et al.[23] conducted a study on colors in context, training classifers to predict colors given multiple colors as reference, demonstrating that color term have a broader meaning when in context with other colors. Butt et al [5] also focus on colors in the context of text-to-image generative models. They propose an approach for optimizing tokens representing target colors. In this work, we study to what extent existing models can depict target colors, considering single and multiple objects and colors. Furthermore, we propose an training-free editing technique which does not require training per RGB color value.
Text-guided Image Editing. The generative power of text-to-image diffusion models has also fueled increasing interests in text-guided image editing [14, 6, 25, 18, 36]. To edit input images, existing methods typically manipulate the internal representations of the denoising networks, in particular the self and cross attention maps [33, 2]. To mitigate color misalignments in text-to-image models, we propose an editing method that builds upon the attention-based framework presented in Liu et al. [18]. We adapt it to our problem setting, as further detailed in Section 4.
@@ 3 The CompColor Benchmark
Despite recent advances in text-to-image generation, it remains unclear whether current models can reliably follow object-level specifications—particularly when prompts involve multiple objects with distinct attributes. For example, when asked to generate an image of “a cornflower-blue gift box with a mint-green ribbon,” we observe that while leading models can correctly render each color when described in isolation, they often fail when multiple colored objects are described together—producing results where the box and ribbon blend into a mixed hue, such as greenish blue, instead of preserving their distinct colors (Figure 1).
@@
Selecting the Color Set. We start by identifying a robust color set that text-to-image models can consistently recognize and generate. This ensures that our benchmark focuses on evaluating compositional ability—rather than testing the model’s ability to understand obscure or rarely used color terms. We begin with the 140 named colors from the HTML color names table [^1] https://en.wikipedia.org/wiki/Web_colors#HTML_color_names. This set includes both basic colors (e.g., Red, Blue) and extended colors (e.g., Lavender, Cyan), providing comprehensive coverage. We evaluate each color using simple prompts (e.g., “a color colored object”) across 5 objects and 10 seeds. Colors consistently recognized by baseline text-to-image models (Stable Diffusion 1.4 and 2.1)—determined by low average color differences in CIELAB space—are retained, resulting in a final set of 35 colors.
@@ 3.1 Evaluation Metrics
We primarily rely on an Accuracy metric, which assigns a score of 1 if the LAB distance between the predicted and target color is below a perceptual threshold $\tau$ (set to 10 empirically), and 0 otherwise. This provides a clear, binary indicator of whether the generated object appears perceptually close to the intended color.
To supplement this with a more fine-grained evaluation, we also compute two additional metrics: (1) LAB Distance, measuring Euclidean distance between the matched cluster color and the ground-truth color in the 3D CIELAB color space, a perceptually uniform color representation, similarly to [1]. (2) RGB L2 Distance Euclidean distance between the matched cluster color and the ground truth color in the RGB space, similarly to [13]. Extended results over these metrics are provided in the supplementary material.
@@ 3.3 Results of Prior Art
Pretrained Models. All pretrained models achieve significantly lower errors when generating single objects, in comparison to multi-object multi-color prompts. This suggests that composing a multi-object image with precise colors is indeed a challenge for image generative models, especially when the prompt specifies distant colors for the different objects. We observe that this is because the generative model tends to switch or leak colors between the objects. For example, when asked to generate a gray colored umbrella and a aqua colored car, it generates an aqua colored umbrella (see the rightmost example on the top row of Figure 4). To further investigate color leakage, we repeated our evaluation under a more lenient criterion: for each object, the generated color is compared against both colors specified in the prompt, and the closest one is used for scoring. This effectively gives the model credit even if colors are swapped or leaked much across objects, thereby isolating and quantifying the impact of color leakage/swap as a common failure mode. Indeed, performance improves significantly (e.g., accuracy of SD 2.1 doubled from 0.26 to 0.5 over distant pairs), demonstrating that leakage is a major factor affecting performance of pretrained models; additional details are provided in the supplementary material.
@@ 4 Method
We aim to address this problem so as to allow users to precisely control the color of the generated objects. For the sake of generality and to allow our approach to be relevant for future generative models as well, we frame our task as an editing task: the input to our system is a prompt $P$ with color attributes (e.g., “A teal cup and a pink spoon”) and a source image $I_{s}$ with the right objects but potentially the wrong colors. This also allows us to edit real images instead of generated ones. To further enable precise control, we assume that the colors come with specified RGB values (e.g., “teal” = (0,128,128) ). Our goal is to edit the image $I_{s}$ so that it faithfully reflects the specified color while preserving its overall appearance and structure and still producing a high quality image.
@@ 4.2 Proposed Approach
We wish to use off-the-shelf text-to-image models to address the task of editing object colors posed above without any further training. This requires us to solve three challenges. First, we want the colors to be matched to the right objects: for example, in the prompt “a teal cup and a pink spoon”, we want the color “pink” to bind to the spoon and not to the cup or the background. Second, we want the color to match the provided RGB values. Third, we want to make these edits while preserving the structure of the image to the extent possible. We address each of these challenges below.
@@ 4.2.1 Binding colors to the right object
The approach taken by SynGen [29] is to add a “contrastive” loss that encourages the attention map of the attribute token (“pink”) to be similar to the attention map of the corresponding object (“spoon”), but very different from the attention map of the other tokens in the prompt. SynGen works fine when the objects are far away and have very distinct attributes, in which case the attention maps are expected to be very distinct. However, when applied to objects that are close and have similar attributes, this approach causes the attributes to be more different than they are supposed to be. This is empirically validated in Section 3 where SynGen actually leads to increased color error when the two objects in the prompt have similar colors.
Instead of a contrastive loss, we propose an alternative approach that constructs pseudo-ground-truth for the cross attention maps. We take inspiration from Stroop probing [3], which is an approach for probing language representations. Stroop probing analyzes language models by comparing the representations of a reference prompt (e.g., ”a photo of a banana”) to a prompt with colors specified (e,g, ”a photo of a blue banana”). Here we take inspiration from this idea and propose to compare the cross-attention maps of our prompt with color (e.g., ”a photo of a blue banana”) with the cross-attention maps of a simplified prompt (” a photo of a banana”). Ideally, we would want the cross-attention maps of ”blue” and ”banana” from the former to match the cross-attention maps of ”banana” from the latter. Thus, the cross-attention maps from the simplified prompt provide pseudo-ground-truth cross-attention maps for all the objects and attributes.
@@ 4.2.2 Getting the right colors
4.2.3 Preserving image structure
Finally, we want to ensure that our modifications do not alter the structure of the original image. To preserve the structural integrity of the original image, we draw inspiration from prior work in diffusion-based image editing [14, 6, 18] and replace the self-attention for the full prompt with the self-attention obtained using the simplified prompt. In addition, to maintain the background, we perform blending at time $t_{b}$ similar to [25] in order to preserve the background appearance during the editing process (see Figure 3). Concretely, for each object, we use the object masks $M^{simp}_{object}$ introduced previously in Section 4.2.2 and take the union to obtain $M^{simp}_{fg}$. Then, we perform blending at time $t_{b}$ during the backward process, updating the latent as follows:
$Z_{t_{b}}\rightarrow Z_{t_{b}}\cdot{M^{simp}_{fg}}+Z^{simp}_{t_{b}}\cdot({1-M^{simp}_{fg}})$ (6)
Putting the ideas described above together, we obtain our proposed approach; see Figure 3. Additional implementation details are provided in the supplementary material.
Method Ours AnySD FPE MasaCtrl
@@ 4.2.3 Preserving image structure
Table 2: Color editing evaluation, comparing our approach to AnySD, FPE and MasaCtrl. We report the absolute improvement in Accuracy(ACC); higher is better, and 0 means no improvement. In green the best absolute improvement in red degradation, results may be sensitive to +-1% due to the color clustering.
5 Evaluation
SD 2.1 Initial        +OursRefer to captionRefer to caption
Initial        +Ours
Refer to captionRefer to caption
Initial        +Ours
Refer to captionRefer to caption
FLUX Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
SynGen Refer to captionRefer to caption ”a hot-pink colored suitcase and a tomato colored hat ” Refer to captionRefer to caption ”a lavender-blush colored flower and a black colored bicycle” Refer to captionRefer to caption ”a gray colored umbrella and a aqua colored car”
@@ 5 Evaluation
Figure 4: Images generated using the baselines, before and after being edited by our approach. We illustrate close color pair in the left example and distant color pairs in the two right examples. Many additional examples are provided in the supplementary material.
In this section, we evaluate our editing approach by comparing it to three editing techniques: AnySD [36], MasaCtrl [6] and FPE [18]. AnySD model introduces task-aware routing and learnable task embedding for unified image editing while the MasaCtrl and FPE manipulate the attention layers for performing text-driven edits. Additional experiments, ablations, and a discussion of limitations are included in the supplementary material. We also provide interactive visualizations over random samples in CompColor, complementing the qualitative results shown in Figure 4, and qualitative comparisons over real images.
As shown in Table 2, our approach significantly improves color accuracy across all methods. These improvements substantially exceed those achieved by existing editing baselines. Moreover, our editing approach yields consistent improvements regardless of the initial image source. This demonstrates that our method is model-agnostic and can generalize to future generative architectures. For both close and distant color pairs, the best results are achieved on images generated by SynGen and edited with our method. Images generated with FLUX also performs well, especially on close color pairs.
Intriguingly, the editing methods sometimes cause a drop in accuracy, particularly for images generated using various versions of Stable Diffusion. This drop can be partially attributed to the difficulty of attention-based methods to bind the colors to the right object in models that often suffer from color swapping, as further discussed in Section 3.3. By contrast, our approach ensures that the right colors are associated with the right object via our Stroop attention loss.

### v2

# Color Bind: Exploring Color Perception in Text-to-Image Models

@@ 1 Introduction
Unfortunately, the answer might be no. Evidence suggests that these models struggle in faithfully generating the semantics conveyed by complex multi-subject prompts [9, 33, 10]. Consider the results depicted in Figure LABEL:fig:teaser. While text-to-image models mostly succeed in generating a “cornflower-blue gift box”, they are challenged by the more complicated prompt that also contains additional colored entities. We observe that a case study on colors offers a unique opportunity for performing a rigorous evaluation over text-to-image generative models. By contrast to coarse CLIP-based metrics, entangled VQA-based scores [20], or small-scale human evaluations conducted by prior work, performance can be measured explicitly using perceptual color differences.
Our contribution is twofold. First, we conduct a set of controlled experiments, comparing results generated using a single object (and its associated color) and results generated using multiple objects (and multiple colors). We also separately consider perceptually-close colors and distant colors, and contribute a benchmark for color understanding in image generation. Second, based on insights from this analysis, we introduce a flexible and model-agnostic approach that addresses these failures not by retraining, but by refining the generated outputs themselves. Concretely, we formulate this as an editing framework that modifies images by accurately binding colors their intended objects while preserving all other attributes (Fig. LABEL:fig:teaser).
Our analysis reveals that pretrained text-to-image models are significantly more successful in representing color attributes in the single subject setting. Inference-time methods that explicitly target the problem of semantic leakage—where attributes intended for one object instead influence unrelated objects—cannot robustly bridge this gap, and also often compromise performance over single subject prompts. We also observe a clear performance gap between close and distant color pairs, demonstrating that in most cases distant color pairs yield worse performance, due to possible leakage which more severely affects performance.
@@ 2 Related Work
Prior work [9, 33] has analyzed and identified typical misalignments, such as catastrophic neglect (i.e., failure to generate one or more mentioned subjects) and incorrect attribute binding (i.e., failure to match between attributes and subjects). Various mitigation schemes have been proposed. These increase semantic alignment by modifying the text embeddings [13, 38] or by better aligning the cross-attention maps [9, 33]. Several techniques mitigating semantic misalignments also utilize additional inputs, such as fine-grained segmentation maps [19], coarse spatial layouts [32, 10] or through the use of compositional generation [23]. In this work, we focus on the problem of incorrect attribute binding, focusing specifically on colors. We introduce a new benchmark for studying such semantic misalignments, analyzing the performance of existing text-to-image models on our proposed benchmark.
@@
Another line of work has explored color understanding for comparing multimodally-trained models with text-only models. Alper et al. [3] demonstrated that vision-and-language models can better associate objects with colors, in comparison to text only models, like BERT [11]. Grimal et al. [15] proposed a new metric for evaluating alignment in text-to-image models, highlighting issues of color leakage and significant performance drops when prompts involve two colors. Several work [28, 22] have studied the effect of reporting bias on language models, demonstrating that they are able to connect object and colors despite the reporting bias in text [26]. Monroe et al.[27] conducted a study on colors in context, training classifers to predict colors given multiple colors as reference, demonstrating that color term have a broader meaning when in context with other colors. Kim et al. [18] noted that diffusion models are less effective at modeling color-related attributes compared to shape-related attributes. Butt et al [6] also focus on colors in the context of text-to-image generative models. They propose an approach for optimizing tokens representing target colors. In this work, we study to what extent existing models can depict target colors, considering single and multiple objects and colors and propose an editing framework which does not require training per RGB color value.
Text-guided Image Editing. The generative power of text-to-image diffusion models has also fueled increasing interests in text-guided image editing [16, 7, 29, 21, 41, 5, 42]. To edit input images, existing methods typically manipulate the internal representations of the denoising networks, in particular the self and cross attention maps [37, 2] or employs instruction-based editing. To mitigate color misalignments in text-to-image models, we propose an editing method that builds upon the attention-based framework presented in Liu et al. [21]. We adapt it to our problem setting, as further detailed in Section 4.
@@ 3 The CompColor Benchmark
Despite recent advances in text-to-image generation, it remains unclear whether current models can reliably follow object-level specifications—particularly when prompts involve multiple objects with distinct attributes. For example, when asked to generate an image of “a cornflower-blue gift box with a mint-green ribbon,” we observe that while leading models can correctly render each color when described in isolation, they often fail when multiple colored objects are described together—producing results where the box and ribbon blend into a mixed hue, such as greenish blue, instead of preserving their distinct colors (Figure LABEL:fig:teaser).
@@
Selecting the Color Set. We start by identifying a robust color set that text-to-image models can consistently recognize and generate. This ensures that our benchmark focuses on evaluating compositional ability—rather than testing the model’s ability to understand obscure or rarely used color terms. We begin with the 140 named colors from the HTML color names table [^1] https://en.wikipedia.org/wiki/Web_colors#HTML_color_names. This set includes both basic colors (e.g., Red, Blue) and extended colors (e.g., Lavender, Cyan), providing comprehensive coverage. We evaluate each color using simple prompts (e.g., “a color colored object”) across 5 objects randomly selected from our object set and 10 seeds. Colors consistently recognized by baseline text-to-image models (Stable Diffusion 1.4 and 2.1)—determined by low average color differences in CIELAB space—are retained, resulting in a final set of 35 colors.
@@ 3.1 Evaluation Metrics
We primarily rely on an Accuracy metric, which assigns a score of 1 if the LAB distance between the predicted and target color is below a perceptual threshold $\tau$, and 0 otherwise. This provides a clear, binary indicator of whether the generated object appears perceptually close to the intended color. In all experiments, we set $\tau=10$; see the supplementary material for additional details and explanations.
To supplement this with a more fine-grained evaluation, we also compute two additional metrics: (1) LAB L2 Distance, measuring Euclidean distance between the matched cluster color and the ground-truth color in the 3D CIELAB color space, a perceptually uniform color representation, similarly to [1]. (2) RGB L2 Distance Euclidean distance between the matched cluster color and the ground truth color in the RGB space, similarly to [14]. Extended results over these metrics are provided in the supplementary.
@@ 3.3 Results of Prior Art
Pretrained Models. All pretrained models achieve significantly lower errors when generating single objects, in comparison to multi-object multi-color prompts. This suggests that composing a multi-object image with precise colors is indeed a challenge for image generative models, especially when the prompt specifies distant colors for the different objects. We observe that this is because the generative model tends to switch or leak colors between the objects. For example, when asked to generate a hot-pink colored suitcase and a tomato colored hat, it generates an tomato colored suitcase (see the leftmost example on the middle row of Figure 4). To further investigate color leakage, we repeated our evaluation under a more lenient criterion: for each object, the generated color is compared against both colors specified in the prompt, and the closest one is used for scoring. This effectively gives the model credit even if colors are swapped or leaked much across objects, thereby isolating and quantifying the impact of color leakage/swap as a common failure mode. Indeed, performance improves significantly (e.g., accuracy of SD 2.1 doubled from 0.26 to 0.5 over distant pairs), demonstrating that leakage is a major factor affecting performance of pretrained models; additional details are provided in the supplementary material.
@@ 4 Method
We aim to address this problem so as to allow users to precisely control the color of the generated objects. For the sake of generality and to allow our approach to be relevant for future generative models as well, we frame our task as an editing task: the input to our system is a prompt $P$ with color attributes (e.g., “A teal cup and a pink spoon”) and a source image $I_{s}$ with the right objects but potentially the wrong colors. This also allows us to edit real images instead of generated ones. To further enable precise control, we assume that the colors come with specified RGB values (e.g., “teal” = (0,128,128) ). Our goal is to edit the image $I_{s}$ so that it faithfully reflects the specified color while preserving its overall appearance and structure and still producing a high quality image; see Figure 3.
@@ 4.2 Proposed Approach
We wish to use off-the-shelf text-to-image models to address the task of editing object colors posed above without any further training. This requires us to solve three challenges. First, we want the colors to be matched to the right objects: for example, in the prompt “a teal cup and a pink spoon”, we want the color “pink” to bind to the spoon and not to the cup or the background. Second, we want the color to match the provided RGB values. Importantly, we want to make these edits while preserving the structure of the image to the extent possible. We address these challenges below.
@@ 4.2.1 Binding colors to the right object
We propose an approach that constructs pseudo-ground-truth for the cross attention maps, different from the contrastive loss proposed by prior work [33] which can degrade performance when objects are close or share similar attributes (as shown in Section 3). We take inspiration from Stroop probing [3], which is an approach for probing language representations. Stroop probing analyzes language models by comparing the representations of a reference prompt (e.g., ”a photo of a banana”) to a prompt with colors specified (e,g, ”a photo of a blue banana”). Here we take inspiration from this idea and propose to compare the cross-attention maps of our prompt with color (e.g., ”a photo of a blue banana”) with the cross-attention maps of a simplified prompt (” a photo of a banana”). Ideally, we would want the cross-attention maps of ”blue” and ”banana” from the former to match the cross-attention maps of ”banana” from the latter. Thus, the cross-attention maps from the simplified prompt provide pseudo-ground-truth cross-attention maps for all the objects and attributes.
@@ 4.2.2 Getting the right colors
Finally, we want to ensure that our modifications do not alter the structure of the original image. To this end, we adopt existing strategies from prior diffusion-based editing methods [16, 7, 21, 29]. Specifically, we replace self-attention maps with those from the simplified prompt and apply background blending during the backward process; additional details are provided in the supplementary.
Method Ours AnySD MagicBrush InstructPix2Pix FPE MasaCtrl
@@
Table 2: Color editing evaluation, comparing our approach to AnySD, MagicBrush, InstructPix2Pix, FPE and MasaCtrl. In addition to the absolute ACC value, we report the improvement (relative to the source image) in colors (green denotes improvements and red denotes degradations).
4.3 Evaluation
SD 2.1 Source Refer to caption +Ours Refer to caption +AnySDRefer to caption +MagicBrushRefer to caption +InstPix2Pix Refer to caption +FPE Refer to caption
+MasaCtrl Refer to caption
FLUX Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
SynGen Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
”a hot-pink colored suitcase and a tomato colored hat”
@@ 4.3 Evaluation
Figure 4: Images generated using the baselines, before and after being edited by our approach, AnySD [41], MagicBrush [42], InstructPix2Pix [5], FPE [21] and MasaCtrl [7]. We illustrate close color pair in the left images. Many additional examples are provided in the supplementary material. Our method consistently outperforms existing editing approaches, we correctly edit the colors and maintain the objects’ shape and background color perfectly aligned with the source image.
To assess the effectiveness of our approach, we compare against several representative editing methods: (i) instruction-following editors (MagicBrush [42], InstructPix2Pix [5], AnySD [41]) trained on paired instruction–image edits, and (ii) attention-manipulation editors (MasaCtrl [7], FPE [21]) that steer self/cross-attention at inference. Implementation details for each method are provided in the supplementary, as well as runtime comparisons, ablations of each module (Stroop loss, color loss, simplified prompt), and results over various metrics that quantify additional aspects, such as semantic alignment and image quality. We also provide interactive visualizations and comparisons over random samples in our benchmark.
Quantitative results. Quantitative results are reported in Table 2. As can be seen in the table, our method consistently improves color accuracy across all base models and settings. The largest improvements are observed when editing SynGen generations; FLUX generations also benefit substantially, especially for close-color prompts. Statistical significance (Wilcoxon signed-rank) and per-model breakdowns are reported in the supplementary.
As can be observed in the table, the improvements obtained by our approach substantially exceed those achieved by existing editing methods. Our experiments demonstrate instruction-following editors are not well designed for multi-color, multi-object binding as they lack a mechanism to explicitly ensure color-object correspondence. In particular, earlier work (i.e., InstructPix2Pix) often fail on this task, producing images that “over-edit” the scene. This is particularly apparent for distant-color pairs, where the model cannot reliably disambiguate which object each color belongs to. Attention-manipulation editors can better preserve the object structure and realism. However, as they don’t directly enforces attribute binding, they frequently mis-bind colors across objects—swapping or blending them—leading to a measurable drop in Accuracy on multi-color prompts. For example, in our SynGen and FLUX comparisons, these methods maintained object layout but assigned the wrong colors to other objects or background regions.
By contrast, our approach introduces two targeted components. The Stroop attention loss explicitly aligns color tokens with the correct object regions by leveraging a color-less pseudo-reference, directly addressing semantic leakage. The color loss further enforces the exact RGB values on the segmented object. Together, these allow our method to outperform baselines across paradigms, delivering edits that are both accurate and structurally consistent.
Qualitative results. Figure 4 shows that our edits faithfully change target colors while preserving object shape and background. For SD 2.1, ours is the only method that performs the intended edit without collateral artifacts. On SynGen, baselines distort object details (e.g., hat shape/ribbon) whereas our edits remain localized and faithful. For FLUX, baselines often tint background regions, while our edits stay confined to the target objects.